# Offline Learning for Combinatorial Multi-armed Bandits

### v1

@@ 2.3 Data Coverage Conditions: Quality of the Dataset
For a CMAB-T instance $\mathcal{I}$ with unknown distribution $\mathbb{D}_{\text{arm}}$ and mean vector $\bm{\mu}$, let $S^{*}=\operatorname*{argmax}_{S\in\mathcal{S}}r(S;\bm{\mu})$.***Note that for simplicity, we choose an arbitrary optimal solution $S^{*}$ , and in practice, we can choose one that leads to the smallest coverage coefficient. We say that the data collecting distribution $\mathbb{D}_{\mathcal{S}}$ satisfies the infinity-norm triggering probability modulated (TPM) data coverage condition, if there exists a coefficient $C^{*}_{\infty}>0$ (referred to as coverage coefficient), we have
@@
For a CMAB-T instance $\mathcal{I}$ with unknown distribution $\mathbb{D}_{\text{arm}}$ and mean vector $\bm{\mu}$, let $S^{*}=\operatorname*{argmax}_{S\in\mathcal{S}}r(S;\bm{\mu})$. We say that the data collecting distribution $\mathbb{D}_{\mathcal{S}}$ satisfies the 1-norm triggering probability modulated (TPM) data coverage condition, if there exists a coefficient $C^{*}_{1}>0$, we have
@@
As will be shown later, our gap upper bound scales with the action size of the optimal action $S^{*}$. Therefore, we give three different metrics to describe the optimal action size, i.e., the number of arms that can be triggered by the optimal action $K^{*}\vcentcolon=\left|\left\{i\in[m]:p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}>0\right\}\right|$, the expected number of arms that can be triggered by the optimal action $\bar{K}^{*}\vcentcolon=\sum_{i\in[m]}p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$, and the $\ell_{2}$ number of arms that can be triggered by the optimal action $\bar{K}_{2}^{*}\vcentcolon=\sum_{i\in[m]}\sqrt{p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}}$. In general, we have $\bar{K}^{*}\leq\bar{K}^{*}_{2}\leq K^{*}$.
@@
Both 3 and 4 evaluate the quality of the dataset $\mathcal{D}$, which directly impacts the amount of data required to accurately estimate the expected reward of the optimal $S^{*}$. The denominator $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}$ represents the data generation rate for arm $i$ and $\frac{1}{p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}}$ corresponds to the expected number of samples needed to observe one instance of arm $i$. Incorporating similar triggering probability modulation as in 2, we use $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ to reweight the importance of each arm $i$, and when $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ is small, the uncertainty associated with arm $i$ has small impact on the estimation. Consequently, a large amount of data is not required for learning about arm $i$. Notably, because we compare against the optimal super arm $S^{*}$, we only require the weight $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ of the optimal action $S^{*}$ as the modulation. This is less restrictive than uniform coverage conditions that require adequate data for all possible actions, as seen in Chen and Jiang [110], Jiang [111].
@@
Note that 3 and 4 are restrictions on the base arm level. Hence, our framework is flexible and can accommodate any data collection distribution $\mathbb{D}_{\mathcal{S}}$, including distributions over actions $\mathcal{S}^{\prime}$ that may assign zero probability to the optimal action $S^{*}$ or even extend beyond the feasible action set $\mathcal{S}$. For example, in LLM cache (Section 4.2), the experimenter might ensure arm feedback by using an empty cache in each round, leveraging cache misses to collect feedback. In this case, the distribution $\mathbb{D}_{\mathcal{S}}$ assigns zero probability to the optimal cache configuration as well as any reasonable cache configurations. See Section 4.2 for further details.
@@ 3 CLCB Algorithm and Theoretical Analysis
The CLCB algorithm first computes high-probability lower confidence bounds (LCBs) for each base arm (line 5). These LCB estimates are then used as inputs to a combinatorial oracle to select an action $\hat{S}$ that approximately maximizes the worst-case reward function $𝝁r(S^{*};\underaccent{\bar}{\bmu})$ (line 7). The key part of Algorithm 1 is to conservatively use the LCB, penalizing each base arm by its confidence interval, $\sqrt{{\log(\frac{4mn}{\delta})}/{2N_{i}}}$.This approach, rooted in the pessimism principle [112], mitigates the impact of high fluctuations in empirical estimates caused by limited observations, effectively addressing the uncertainty inherent in passively collected data.
@@
The proof of Theorem 1 consists of three key steps: (1) express the suboptimality gap in terms of the uncertainty gap $𝝁r(S^{*};\bm{\mu})-r\left(S^{*};\underaccent{\bar}{\bmu}\right)$ over the optimal action $S^{*}$, rather than the on-policy error over the chosen action $\hat{S}$ as in online CMAB, (2) leverage 2 to relate the uncertainty gap to the per-arm estimation gap, and (3) utilize 2 to deal with the arbitrary data collection probabilities and bound the per-arm estimation gap in terms of $n$. For a detailed proof, see Appendix A. ∎
@@
Looking at the suboptimality gap result, both Eq. 5 and Eq. 6 decrease at a rate of $\frac{1}{\sqrt{n}}$ with respect to the number of offline data samples $n$. Additionally, they scale linearly with the smoothness coefficient $B_{1}$ and the approximation ratio $\alpha$. For problems satisfying Eq. 5, the gap scales linearly with the $\ell_{2}$-action size $\bar{K}_{2}^{*}$ and the the coverage coefficient $C_{\infty}^{*}$ in Eq. 5. For problems satisfying Eq. 6, the gap depends on the action size $\bar{K}^{*}$ and the 1-norm data coverage coefficient $C_{1}^{*}$. To output an action that is $\epsilon$-close to $S^{*}$, Eq. 5 and Eq. 6 need $\tilde{O}(B_{1}^{2}\alpha^{2}{\bar{K}_{2}^{*}}^{2}C_{\infty}^{*}/\epsilon^{2})$ and $\tilde{O}(B_{1}^{2}\alpha^{2}{\bar{K}^{*}}C_{1}^{*}/\epsilon^{2})$ samples, respectively. In general, we have $C_{1}^{*}\leq K^{*}C_{\infty}^{*}$ and $\bar{K}^{*}\geq\frac{(\bar{K}^{*}_{2})^{2}}{K^{*}}$, indicating that neither Eq. 5 nor Eq. 6 strictly dominates the other. For instance, for CMAB with semi-bandit feedback where $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}=p_{j}^{\mathbb{D}_{\text{arm}},S^{*}}=1$ for any $i,j\in S^{*}$ and 00 otherwise, Eq. 6 is tighter than Eq. 5 since $\bar{K}^{*}=\bar{K}^{*}_{2}=K^{*}$ and $C_{1}^{*}\leq K^{*}C^{*}_{\infty}$. Conversely, for the LLM cache to be introduced in Section 4.2, if the experimenter selects the empty cache each time, such that $\frac{p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}}{p_{i}^{\mathbb{D}_{s},S^{*}}}=1$ for $i\in S^{*}$, then we have $C_{1}^{*}=K^{*}C^{*}_{\infty}$. Since $\bar{K}^{*}\geq\frac{(\bar{K}^{*}_{2})^{2}}{K^{*}}$ so Eq. 5 is tighter than Eq. 6.
Lower bound result. In this section, we establish the lower bound for a specific combinatorial multi-armed bandit (CMAB) problem: the stochastic $k$-path problem $\mathcal{I}$. This problem was first introduced in [6] to derive lower bounds for the online CMAB problem.
@@ 4 Applications of the Off-CMAB Framework
In this section, we introduce three representative applications that can fit into our Off-CMAB framework with new/improved results, which are summarized in Table 1. We also provide empirical evaluations for the cascading bandit and the LLM cache in Section 5.
@@
LLM Cache (Section 4.2) $B_{1}=1$ $C_{1}^{*}=m$ $\tilde{O}\left(\sqrt{\frac{m}{n}}\right)$ $\tilde{O}\left(\sqrt{\frac{k^{2}}{C_{1}}}\right)^{\dagger}$ over Zhu et al. [34]
Social Influence Maximization (Section 4.3) $B_{1}=V$ 1∗∗ $\tilde{O}\left(\sqrt{\frac{V^{2}d^{2}_{\max}\sigma^{2}(S^{*};G)}{\eta\cdot\gamma^{3}\cdot n}}\right)$ $\tilde{O}\left(\sqrt{\frac{V^{4}}{k^{2}d^{2}_{\max}\eta}}\right)^{\ddagger}$ over Chen et al. [107]
@@
  • ∗∗ Similar to Chen et al. [107], we depend directly on assumption for seed sampling probability bound $\gamma$ and the activation probability bound $\eta$;
@@ 4.1 Offline Learning for Cascading Bandits
Learning to rank [33] is an approach used to improve the ordering of items (e.g., products, ads) in recommendation systems based on user interactions. This approach is crucial for various types of recommendation systems, such as search engines [113], e-commerce platforms [114], and content recommendation services [115].
The cascading bandit problem [18, 19, 79] addresses the online learning to rank problem under the cascade model [116]. The canonical cascading bandit problem considers a $T$-round sequential decision-making process. At each round $t\in[T]$, a user $t$ comes to the recommendation system (e.g., Amazon), and the learner aims to recommend a ranked list $S_{t}=(a_{t,1},...,a_{t,k})\subseteq[m]$ of length $k$ (i.e., a super arm) from a total of $m$ candidate products (i.e., base arms). Each item $i\in S_{t}$ has an unknown probability $\mu_{i}$ of being satisfactory and purchased by user $t$, which without loss of generality, is assumed to be in descending order $\mu_{1}\geq\mu_{2}\geq...\geq\mu_{m}$.
@@
2:for arm $q\in{\mathcal{Q}}$ do
6:end for
8:Return: $\hat{\mathcal{M}}$.
Algorithm and result. This application fits into the CMAB-T framework, satisfying 2 with coefficient $B_{1}=1$ as in [9]. The oracle is essentially to find the top-$k$ items regarding LCB $\underaccent{\bar}{\mu}_{i}$, which maximizes $𝝁r(S;\underaccent{\bar}{\bmu})$ in $O(m\log k)$ time complexity using the max-heap. Plugging this oracle into line 7 of Algorithm 1 gives the algorithm, whose detail is in Algorithm 5 in Appendix C.
@@
$r(S^{*};\bm{\mu})-r\left(\hat{S};\bm{\mu}\right)\leq 2\sqrt{\frac{2k\log(\frac{2mn}{\delta})}{n}\sum_{i=1}^{k}\frac{\prod_{j=1}^{i-1}(1-\mu_{j})}{\sum_{j=1}^{k}q_{ij}(1-\mu_{1})^{j-1}}},$ (7)
If $\,\mathbb{D}_{\mathcal{S}}$ is a uniform distribution so that $q_{ij}=\frac{1}{m}$, it holds that
@@ 4.2 Offline Learning for LLM Cache
The LLM cache is a system designed to store and retrieve outputs of Large Language Models (LLMs), aiming to enhance efficiency and reduce redundant computations during inference [91]. As LLMs are increasingly integrated into real-world applications, including conversational AI, code generation, and personalized recommendations, efficient caching strategies are critical to improving the scalability, responsiveness, and resource utilization of these systems, particularly in large-scale, real-time deployments [88, 91, 34].
@@
In the LLM cache bandit [34], which is a $T$-round sequential learning problem, we consider a finite set of $m$ distinct queries ${\mathcal{Q}}=\{q_{1},...,q_{m}\}$. Each query $q\in{\mathcal{Q}}$ is associated with an unknown expected cost $c(q)\in[0,1]$ and unknown probability $p(q)\in[0,1]$. From the LLM cache bandit point of view, we have $2m$ base arms: the first $m$ arms correspond to the unknown costs $c(q)\in[0,1]$ for $q\in{\mathcal{Q}}$, and the last $m$ arms correspond to the probability $p(q)\in[0,1]$ for $q\in{\mathcal{Q}}$. When query $q$ is input to the LLM system, the LLM processes it and returns a corresponding response (i.e., answer) $r(q)$. Every round $t$ when the LLM processes $q$, it will incur a random cost $C_{t}(q)$ with mean $c(q)$, where $C_{t}(q)=c(q)+\epsilon_{t}(q)$, where $\epsilon_{t}(q)$ is a sub-Gaussian noise that captures the uncertainties in the cost, with $\mathbb{E}[\epsilon_{t}(q)]=0$. The cost in an LLM can be floating point operations (FLOPs), latency of the model, the price for API calls, user satisfaction of the results, or a combination of all these factors. The goal of LLM cache bandit is to find the optimal cache $\mathcal{M}^{*}$ storing the query-response pairs that are both likely to be reused and associated with high costs.
@@
Expected cost function and cache feedback. In each round $t$, a user comes to the system with query $q_{t}$, which is sampled from $\mathcal{Q}$ according to a fixed unknown distribution $\{p(q)\}_{q\in{\mathcal{Q}}}$ with $\sum_{q\in{\mathcal{Q}}}p(q)=1$. To save the cost of repeatedly processing the queries, the LLM system maintains a cache $\mathcal{M}_{t}\subseteq{\mathcal{Q}}$, storing a small subset of queries of size $k\geq 0$ with their corresponding results. After $q_{t}$ is sampled, the agent will first check the current cache $\mathcal{M}_{t}$. If the query $q_{t}$ is found in the cache, i.e., $q_{t}\in\mathcal{M}_{t}$, we say the query hits the cache. In this case, the result of $q_{t}$ is directly returned without further processing by the LLM. The cost of processing this query is 00 and will save a potential cost $C_{t}(q_{t})$, which is unobserved to the agent. If query $q_{t}$ does not hit the cache, the system processes the query, incurring a cost $C_{t}(q_{t})$ which is observed by the learner, and returns the result $r(q_{t})$. Let us denote $\bm{c}=(c(q))_{q\in{\mathcal{Q}}}$ and $\bm{p}=(p(q))_{q\in{\mathcal{Q}}}$ for convenience. Given any cache $\mathcal{M}$ and the query $q$, the random cost saved in round $t$ is $C_{t}(\mathcal{M},q)=\mathbb{I}\{q\notin\mathcal{M}\}C_{t}(q)$. Thus the expected cost incurred is $c(\mathcal{M};\bm{c},\bm{p})=\mathbb{E}\left[\sum_{q\in{\mathcal{Q}}}C_{t}(\mathcal{M},q)\right]=\sum_{q\notin\mathcal{M}}p(q)c(q)$.
@@
Algorithm and result. This application fits into the CMAB-T framework, satisfying the 1-norm TPM smoothness condition with coefficient $B_{1}=1$ (see Section D.1 for the detailed proof). Since we are minimizing the cost rather than maximizing the reward, we use the UCB $\bar{p}(q)\bar{c}(q)$. The oracle is essentially to find the top-$k$ queries regarding UCB $\bar{p}(q)\bar{c}(q)$, which minimizes $c(S;\bar{\bm{c}},\bar{\bm{p}})$ in $O(m\log k)$ time complexity using the max-heap. The detailed algorithm and its result is provided in Algorithm 2.
@@
$c(\mathcal{M}^{*};\bm{c},\bm{p})-c\left(\hat{\mathcal{M}};\bm{c},\bm{p}\right)\leq 2\sqrt{\frac{2\sum_{q\in{\mathcal{Q}}-\mathcal{M}^{*}}{\frac{1}{\nu(q)}}\log(\frac{6mn}{\delta})}{n}}+2\sqrt{\frac{2m\log(\frac{3}{\delta})}{n}},$ (9)
If the experimenter samples empty cache $\mathcal{M}_{t}=\emptyset$ in each round as in [34] so that $\nu(q)=1$, it holds that
@@
Looking at Eq. 10, our result improves upon the state-of-the-art result [34] $\tilde{O}(k\sqrt{\frac{m}{C_{1}n}})$ by a factor of $\tilde{O}(\sqrt{\frac{k^{2}}{C_{1}}})$, where $C_{1}>0$ is assumed to be an lower bound of $c(q)$ for $q\in{\mathcal{Q}}$. This improvement comes from our tight analysis to deal with the triggering probability (the $1/C_{1}$ can be thought of as minimum triggering probability in their analysis) and from the way that we deal with full-feedback arm $\bm{p}$. Furthermore, since we use the UCB $\bar{\bm{c}}$ while they use LCB $𝐜\underaccent{\bar}{\bc}$, our algorithm in principle can be generalized to more complex distributions as long as the optimal queries are sufficiently covered. As a by-product, we also consider the online streaming LLM cache bandit setting as in [34] from our CMAB-T point of view, which improves their result $\tilde{O}(\frac{km\sqrt{T}}{C_{1}})$ by a factor of $\tilde{O}(\frac{k\sqrt{m}}{C_{1}})$. We defer the details to Section D.3.

### v2

@@ 2.3 Data Coverage Conditions: Quality of the Dataset
For a CMAB-T instance $\mathcal{I}$ with unknown distribution $\mathbb{D}_{\text{arm}}$ and mean vector $\bm{\mu}$, let $S^{*}=\operatorname*{argmax}_{S\in\mathcal{S}}r(S;\bm{\mu})$. [^2] Note that for simplicity, we choose an arbitrary optimal solution $S^{*}$ , and in practice, we can choose one that leads to the smallest coverage coefficient. We say that the data collecting distribution $\mathbb{D}_{\mathcal{S}}$ satisfies the infinity-norm triggering probability modulated (TPM) data coverage condition, if there exists a coefficient $C^{*}_{\infty}>0$ (referred to as coverage coefficient), such that
@@
For a CMAB-T instance $\mathcal{I}$ with unknown distribution $\mathbb{D}_{\text{arm}}$ and mean vector $\bm{\mu}$, let $S^{*}=\operatorname*{argmax}_{S\in\mathcal{S}}r(S;\bm{\mu})$. We say that the data collecting distribution $\mathbb{D}_{\mathcal{S}}$ satisfies the 1-norm triggering probability modulated (TPM) data coverage condition, if there exists a coefficient $C^{*}_{1}>0$, such that
@@
Both 3 and 4 evaluate the quality of the dataset $\mathcal{D}$, which directly impacts the amount of data required to accurately estimate the expected reward of the optimal $S^{*}$. The denominator $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}$ represents the data generation rate for arm $i$, and $\frac{1}{p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}}$ corresponds to the expected number of samples needed to observe one instance of arm $i$. Incorporating similar triggering probability modulation as in 2, we use $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ to reweight the importance of each arm $i$, and when $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ is small, the uncertainty associated with arm $i$ has small impact on the estimation. Consequently, a large amount of data is not required for learning about arm $i$. Notably, because we compare against the optimal super arm $S^{*}$, we only require the weight $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}$ of the optimal action $S^{*}$ as the modulation. This is less restrictive than uniform coverage conditions that require adequate data for all possible actions, as used by Chen & Jiang (2019b); Jiang (2019).
@@
Note that 3 and 4 are restrictions on the base arm level. Hence, our framework is flexible and can accommodate any data collection distribution $\mathbb{D}_{\mathcal{S}}$, including distributions over actions $\mathcal{S}^{\prime}$ that may assign zero probability to the optimal action $S^{*}$ or even extend beyond the feasible action set $\mathcal{S}$. For example, in the LLM cache problem (Section 4.2), the experimenter might ensure arm feedback by using an empty cache in each round, leveraging cache misses to collect feedback. In this case, the distribution $\mathbb{D}_{\mathcal{S}}$ assigns zero probability to the optimal cache configuration as well as any reasonable cache configurations.
@@ 3 CLCB Algorithm and Theoretical Analysis
The CLCB algorithm first computes high-probability lower confidence bounds (LCBs) for each base arm (line 5). These LCB estimates are then used as inputs to a combinatorial oracle to select an action $\hat{S}$ that approximately maximizes the worst-case reward function $𝝁r(S^{*};\underaccent{\bar}{\bmu})$ (line 7). The key part of Algorithm 1 is to conservatively use the LCB, penalizing each base arm by its confidence interval, $\sqrt{{\log(\frac{4mn}{\delta})}/{2N_{i}}}$.This approach, rooted in the pessimism principle (Jin et al., 2020a), mitigates the impact of high fluctuations in empirical estimates caused by limited observations, effectively addressing the uncertainty inherent in passively collected data.
@@
The proof of Theorem 1 consists of three key steps: (1) express the suboptimality gap in terms of the uncertainty gap $𝝁r(S^{*};\bm{\mu})-r\left(S^{*};\underaccent{\bar}{\bmu}\right)$ over the optimal action $S^{*}$, rather than the on-policy error over the chosen action $\hat{S}$ as in online CMAB, (2) leverage 2 to relate the uncertainty gap to the per-arm estimation gap, and (3) utilize 2 to deal with the arbitrary data collection probabilities and bound the per-arm estimation gap in terms of $n$. For a detailed proof, see Appendix D. ∎
@@
Looking at the suboptimality gap result, both Eq. 5 and Eq. 6 decrease at a rate of $\frac{1}{\sqrt{n}}$ with respect to the number of offline data samples $n$. Additionally, they scale linearly with the smoothness coefficient $B_{1}$ and the approximation ratio $\alpha$. For problems satisfying Eq. 5, the gap scales linearly with the $\ell_{2}$-action size $\bar{K}_{2}^{*}$ and the the coverage coefficient $C_{\infty}^{*}$ in Eq. 5. For problems satisfying Eq. 6, the gap depends on the action size $\bar{K}^{*}$ and the 1-norm data coverage coefficient $C_{1}^{*}$. To output an action that is $\epsilon$-close to $S^{*}$, Eq. 5 and Eq. 6 need $\tilde{O}(B_{1}^{2}\alpha^{2}{\bar{K}_{2}^{*}}^{2}C_{\infty}^{*}/\epsilon^{2})$ and $\tilde{O}(B_{1}^{2}\alpha^{2}{\bar{K}^{*}}C_{1}^{*}/\epsilon^{2})$ samples, respectively.
In general, we have $C_{1}^{*}\leq K^{*}C_{\infty}^{*}$ and $\bar{K}^{*}\geq\frac{(\bar{K}^{*}_{2})^{2}}{K^{*}}$, indicating that neither Eq. 5 nor Eq. 6 strictly dominates the other. For instance, for CMAB with semi-bandit feedback where $p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}=p_{j}^{\mathbb{D}_{\text{arm}},S^{*}}=1$ for any $i,j\in S^{*}$ and 00 otherwise, Eq. 6 is tighter than Eq. 5 since $\bar{K}^{*}=\bar{K}^{*}_{2}=K^{*}$ and $C_{1}^{*}\leq K^{*}C^{*}_{\infty}$. Conversely, for the LLM cache to be introduced in Section 4.2, if the experimenter selects the empty cache each time, such that $\frac{p_{i}^{\mathbb{D}_{\text{arm}},S^{*}}}{p_{i}^{\mathbb{D}_{s},S^{*}}}=1$ for $i\in S^{*}$, then we have $C_{1}^{*}=K^{*}C^{*}_{\infty}$. Since $\bar{K}^{*}\geq\frac{(\bar{K}^{*}_{2})^{2}}{K^{*}}$ so Eq. 5 is tighter than Eq. 6.
Lower bound result. In this section, we establish the lower bound for a specific combinatorial multi-armed bandit (CMAB) problem: the stochastic $k$-path problem $\mathcal{I}$. This problem was first introduced in (Kveton et al., 2015c) to derive lower bounds for the online CMAB problem.
@@ 4 Applications of the Off-CMAB Framework
In this section, we introduce three representative applications that can fit into our Off-CMAB-T framework with new/improved results, which are summarized in Table 1. We also provide empirical evaluations for the cascading bandit and the LLM cache in Section 5.
@@
LLM Cache (Section 4.2) $B_{1}=1$ $C_{1}^{*}=m$ $\tilde{O}\left(\sqrt{\frac{m}{n}}\right)$ $\tilde{O}\left(\sqrt{\frac{k^{2}}{C_{1}}}\right)^{\dagger}$
Social Influence Maximization (Section 4.3) $B_{1}=V$ 1∗∗ $\tilde{O}\left(\sqrt{\frac{V^{2}d^{2}_{\max}\sigma^{2}(S^{*};G)}{\eta\cdot\gamma^{3}\cdot n}}\right)$ $\tilde{O}\left(\sqrt{\frac{V^{4}}{k^{2}d^{2}_{\max}\eta}}\right)^{\ddagger}$
@@
  • ∗∗ Similar to Chen et al. (2021), we depend directly on assumption for seed sampling probability bound $\gamma$ and the activation probability bound $\eta$;
@@ 4.1 Offline Learning for Cascading Bandits
The cascading bandit problem (Kveton et al., 2015a; Li et al., 2016; Vial et al., 2022; Dai et al., 2025a) addresses the online learning to rank problem (Liu et al., 2009) under the cascade model (Craswell et al., 2008). The canonical cascading bandit problem considers a $T$-round sequential decision-making process. At each round $t\in[T]$, a user $t$ comes to the recommendation system (e.g., Amazon), and the learner aims to recommend a ranked list $S_{t}=(a_{t,1},...,a_{t,k})\subseteq[m]$ of length $k$ (i.e., a super arm) from a total of $m$ candidate products (i.e., base arms). Each item $i\in S_{t}$ has an unknown probability $\mu_{i}$ of being satisfactory and purchased by user $t$, which without loss of generality, is assumed to be in descending order $\mu_{1}\geq\mu_{2}\geq...\geq\mu_{m}$.
@@
Algorithm and result. This application fits into the CMAB-T framework, satisfying 2 with coefficient $B_{1}=1$ as in (Wang & Chen, 2017). The oracle is essentially to find the top-$k$ items regarding LCB $\underaccent{\bar}{\mu}_{i}$, which maximizes $𝝁r(S;\underaccent{\bar}{\bmu})$ in $O(m\log k)$ time complexity using the max-heap. Plugging this oracle into line 7 of Algorithm 1 gives the algorithm, whose detail is in Algorithm 4 in Appendix F.
@@
$r(S^{*};\bm{\mu})-r\left(\hat{S};\bm{\mu}\right)$
$\leq 2\sqrt{\frac{2k\log(\frac{2mn}{\delta})}{n}\sum_{i=1}^{k}\frac{\prod_{j=1}^{i-1}(1-\mu_{j})}{\sum_{j=1}^{k}q_{ij}(1-\mu_{1})^{j-1}}},$ (7)
If $\,\mathbb{D}_{\mathcal{S}}$ is a uniform distribution so that $q_{ij}=\frac{1}{m}$, it holds that $C_{1}^{*}\leq\frac{\mu_{1}\cdot m}{\mu_{k}}$ and
@@ 4.2 Offline Learning for LLM Cache
The LLM cache is a system designed to store and retrieve outputs of Large Language Models (LLMs), aiming to enhance efficiency and reduce redundant computations during inference (Pope et al., 2022; Bang, 2023; Zhu et al., 2023; Dai et al., 2025b).
@@
In the LLM cache bandit (Zhu et al., 2023), which is a $T$-round sequential learning problem, we consider a finite set of $m$ distinct queries ${\mathcal{Q}}=\{q_{1},...,q_{m}\}$. Each query $q\in{\mathcal{Q}}$ is associated with an unknown expected cost $c(q)\in[0,1]$ and unknown probability $p(q)\in[0,1]$, for a total of $2m$ base arms. When query $q$ is input to the LLM system, the LLM processes it and returns a corresponding response (i.e., answer) $r(q)$. Every round $t$ when the LLM processes $q$, it will incur a random cost $C_{t}(q)$ with mean $c(q)$, representing floating point operations (FLOPs) or the price for API calls (Zhu et al., 2023). We assume that $C_{t}(q)=c(q)+\epsilon_{t}(q)$, where $\epsilon_{t}(q)$ is a sub-Gaussian noise that captures the uncertainties in the cost, with $\mathbb{E}[\epsilon_{t}(q)]=0$.
The goal of LLM cache bandit is to find the optimal cache $\mathcal{M}^{*}$ storing the query-response pairs that are both likely to be reused and associated with high costs.
@@
Expected cost function and cache feedback. In each round $t$, a user comes to the system with query $q_{t}$, which is sampled from $\mathcal{Q}$ according to a fixed unknown distribution $\{p(q)\}_{q\in{\mathcal{Q}}}$ with $\sum_{q\in{\mathcal{Q}}}p(q)=1$. To save the cost of repeatedly processing the queries, the LLM system maintains a cache $\mathcal{M}_{t}\subseteq{\mathcal{Q}}$, storing a small subset of queries of size $k\geq 0$ with their corresponding results. After $q_{t}$ is sampled, the agent will first check the current cache $\mathcal{M}_{t}$. If the query $q_{t}$ is found in the cache, i.e., $q_{t}\in\mathcal{M}_{t}$, we say the query hits the cache. In this case, the result of $q_{t}$ is directly returned without further processing by the LLM. The cost of processing this query is 00 and will save a potential cost $C_{t}(q_{t})$, which is unobserved to the agent. If query $q_{t}$ does not hit the cache, the system processes the query, incurring a cost $C_{t}(q_{t})$ which is observed by the learner, and returns the result $r(q_{t})$. Let us denote $\bm{c}=(c(q))_{q\in{\mathcal{Q}}}$ and $\bm{p}=(p(q))_{q\in{\mathcal{Q}}}$ for convenience. Given any cache $\mathcal{M}$ and the query $q$, the random cost saved in round $t$ is $C_{t}(\mathcal{M},q)=\mathbb{I}\{q\notin\mathcal{M}\}C_{t}(q)$. Thus the expected cost incurred is
$c(\mathcal{M};\bm{c},\bm{p})=\mathbb{E}\left[\sum_{q\in{\mathcal{Q}}}C_{t}(\mathcal{M},q)\right]=\sum_{q\notin\mathcal{M}}p(q)c(q).$ (9)
@@
Algorithm and result. This application fits into the CMAB-T framework, satisfying the 1-norm TPM smoothness condition with coefficient $B_{1}=1$ (see Section G.1 for the detailed proof). Since we are minimizing the cost rather than maximizing the reward, we use the UCB $\bar{p}(q)\bar{c}(q)$. The oracle is essentially to find the top-$k$ queries regarding UCB $\bar{p}(q)\bar{c}(q)$, which minimizes $c(S;\bar{\bm{c}},\bar{\bm{p}})$ in $O(m\log k)$ time complexity using the max-heap. The detailed algorithm and its result is provided in Algorithm 5.
@@
$c(\mathcal{M}^{*};\bm{c},\bm{p})-c\left(\hat{\mathcal{M}};\bm{c},\bm{p}\right)$ (10)
$\leq 2\sqrt{\frac{2\sum_{q\in{\mathcal{Q}}-\mathcal{M}^{*}}{\frac{1}{\nu(q)}}\log(\frac{6mn}{\delta})}{n}}+2\sqrt{\frac{2m\log(\frac{3}{\delta})}{n}},$
If the experimenter samples empty cache $\mathcal{M}_{t}=\emptyset$ in each round as in (Zhu et al., 2023) so that $\nu(q)=1$, it holds that
@@
Looking at Eq. 11, our result improves upon the state-of-the-art result (Zhu et al., 2023) $\tilde{O}(k\sqrt{\frac{m}{C_{1}n}})$ by a factor of $\tilde{O}(\sqrt{\frac{k^{2}}{C_{1}}})$, where $C_{1}>0$ is assumed to be an lower bound of $c(q)$ for $q\in{\mathcal{Q}}$. This improvement comes from our tight analysis to deal with the triggering probability (the $1/C_{1}$ can be thought of as minimum triggering probability in their analysis) and from the way that we deal with full-feedback arm $\bm{p}$. Furthermore, since we use the UCB $\bar{\bm{c}}$ while they use LCB $𝐜\underaccent{\bar}{\bc}$, our algorithm in principle can be generalized to more complex distributions as long as the optimal queries are sufficiently covered. As a by-product, we also consider the online streaming LLM cache bandit setting as in (Zhu et al., 2023) from our CMAB-T point of view, which improves their result $\tilde{O}(\frac{km\sqrt{T}}{C_{1}})$ by a factor of $\tilde{O}(\frac{k\sqrt{m}}{C_{1}})$. We defer the details to Section G.3.